# “Haet Bhasha aur Diskrimineshun”: Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs

### v3

@@ Abstract
Recently released LLMs have strong multilingual & multimodal capabilities. Model vulnerabilities are exposed using audits and red-teaming efforts. Existing efforts have focused primarily on the English language; thus, models continue to be susceptible to multilingual jailbreaking strategies, especially for multimodal contexts. In this study, we introduce a novel strategy that leverages code-mixing and phonetic perturbations to jailbreak LLMs for both text and image generation tasks. We also present an extension to a current jailbreak-template-based strategy and propose a novel template, showing higher effectiveness than baselines. Our work presents a method to effectively bypass safety filters in LLMs while maintaining interpretability by applying phonetic misspellings to sensitive words in code-mixed prompts. We achieve a 99% Attack Success Rate for text generation and 78% for image generation, with Attack Relevance Rate of 100% for text generation and 96% for image generation for the phonetically perturbed code-mixed prompts. Our interpretability experiments reveal that phonetic perturbations impact word tokenization, leading to jailbreak success. Our study motivates increasing the focus towards more generalizable safety alignment for multilingual multimodal models, especially in real-world settings wherein prompts can have misspelt words.
@@
“Haet Bhasha aur Diskrimineshun”:
Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs
@@ 1 Introduction
Large language models (LLMs) are used for a variety of general-purpose Hadi et al. (2023) and safety-critical Hua et al. (2024) tasks by a diverse set of users worldwide. These models are widely accessible via web-based chat interfaces and economically priced APIs [^1] https://openai.com/api/pricing/, and their growing usage has led to increased scrutiny, with a large focus on safety Salhab et al. (2024), bias & hallucination Lin et al. (2024) and privacy violations Das et al. (2025). Red teaming Sarkar (2025), a key evaluation method, uses novel prompting strategies Pang et al. (2025) to bypass safety filters of LLMs and elicit harmful or unethical responses Wei et al. (2023) that exposes model biases and shortcomings.
Code-mixing (CM): Code mixing, the practice of mixing multiple languages within the same conversation, has significantly helped improve the multilingual performance of models for various NLP tasks such as Sentiment Analysis Lal et al. (2019), Machine Translation Chatterjee et al. (2023) and Hate Speech Detection Bohra et al. (2018). With this improved LLM multilingualism, newer and harder bias Mihaylov and Shtedritski (2024) & alignment Shen et al. (2024a) challenges have cropped up, which need to be addressed to ensure fair and safe performance.
Phonetic perturbations: Perturbations are small, intentional changes at various stages of a model pipeline to evaluate the robustness of a network. Multiple techniques have been proposed for both vision Akhtar et al. (2021); Chakraborty et al. (2021); Hendrycks and Dietterich (2019); Wang et al. (2024) and text Goyal et al. (2023); Romero-Alvarado et al. (2024); Moradi and Samwald (2021) domains in the literature. Phonetic perturbations alter a word’s spelling while preserving its pronunciation and semantic meaning.
In societies where English is not the first language, non-native speakers often create new, possibly strange spellings for words based on their phonetic perceptions. For example, ‘design’ and ‘dezain’ have very different spellings but the same pronunciation. While the former is a regular word in the English dictionary, the latter has no meaning. This is often observed in textese Drouin (2011), a form of communication common in SMS and internet conversations Thakur (2021). Such textese-generated spellings manifest as inadvertent perturbations when LLMs are used by non-native speakers. In this work, we present a novel jailbreak technique that combines code-mixing (CM) using the Hindi language with phonetic perturbations (CMP) – injecting phonetically similar sounding spelling errors in sensitive words – to make the challenge more generalizable for a real-world setting. Our study poses the following research questions.
@@
RQ1. Do the safety guardrails of SOTA LLMs generalise to textese-style code-mixed inputs? With their proliferation, LLMs face increasingly novel scenarios like non-native speakers [^2] https://www.bcg.com/publications/2024/consumers-know-more-about-ai-than-businesses-think naturally using their established, informal styles on these new platforms. While models undergo extensive alignment stages Dubey et al. (2024), it is crucial to evaluate if these safety measures hold up in such cases. Our results indicate that our novel strategy results in significant boosts in harmful outputs, revealing the lack of generalisability of existing safeguards.
RQ2. Does our red-teaming attack trigger relevant responses from the LLMs? Phonetic perturbations intentionally misspell words while preserving the pronunciation. Thus, it is necessary to determine whether models correctly interpret the input prompt or if they generate irrelevant content. We demonstrate that LLMs generate harmful content that is highly relevant to the original prompt, indicating failure of safety filters despite high interpretability.
@@
RQ3. How do the phonetic perturbations successfully bypass the safety filters of LLMs? Beyond the success of the attack, it is important to identify the actual vulnerability behind this safety failure. To this end, we employ the method of Integrated Gradients Sundararajan et al. (2017) to conduct an interpretability analysis on one of the open-source LLMs, revealing that phonetic perturbations alter word tokenization in a way that prevents safety filters from triggering.
RQ4. Do phonetic perturbations work successfully for red-teaming large image generation models? The burgeoning interest in extending LLMs to multiple modalities has increased the scope of jailbreak attacks, such as generating fake Ellery (2023) or harmful content Ghosh et al. (2024), making it crucial to evaluate model capabilities across modalities. We demonstrate that our approach generalizes to image generation, producing highly harmful outputs with Multimodal LLMs (MLLMs).
@@ 2 Related Work
Red teaming Ganguli et al. (2022) tasks focus on evaluating LLMs for safety and vulnerability concerns Bhardwaj and Poria (2023). Jailbreaking is one such method that involves bypassing the safety training of LLMs to elicit harmful, unethical, or unintended outputs. Red-teaming tasks on LLMs Chen et al. (2025); Liu et al. (2023) and VLMs Liu et al. (2024) study various forms of vulnerabilities.
Code-mixing (CM) increases the performance Shankar et al. (2024) and capabilities Zhang et al. (2024) of LLMs in multilingual settings. Prior works in multilingual red-teaming of LLMs involves evaluation Shen et al. (2024a), jailbreaking Deng et al. (2023) as well as alignment Song et al. (2024) strategies. Unlike code-switching Yoo et al. (2024) which directly inserts words from other languages in their native scripts, code-mixing uses the script of the primary language to insert words from other languages, making it a more natural communication style as observed in SMS and internet conversations Thakur (2021). In this study, we study code-mixing with phonetic perturbations of sensitive words as an attack vector to evaluate safety alignment of LLMs. This novel jailbreak strategy successfully jailbreaks even SOTA models like Llama 3 and ChatGPT 4o-mini.
3 Datasets, Models & Jailbreaks
We first describe the benchmark datasets and the models evaluated as part of our red-teaming task, followed by a brief description of the standard jailbreak templates and our own proposed templates.
3.1 Datasets Benchmarked
Prompts for text generation: We utilize three benchmark datasets 3HarmfulQA Bhardwaj and Poria (2023), NicheHazardQA Hazra et al. (2024) and TechHazardQA Banerjee et al. (2024) – studying model vulnerabilities, refusal training and compliance with harmful queries. For a comprehensive evaluation of our red-teaming strategy, we sample 20 prompts from each category in each dataset, yielding a total of 460 prompts across 23 categories. All prompts are originally in the English language. We attempted to automate Le et al. (2022) the generation of CMP prompts, but the results were unsatisfactory. Thus, using the sampled datasets, we manually generate the CM and CMP prompt sets, described later.
Prompts for image generation: Using a set of 10 handwritten samples, we prompt GPT-4o to automatically generate sets of 20 red-teaming prompts each to test the model’s resilience against various categories of harm – Religious Hate, Casteist Hate, Gore, Self-Harm and Social Media Toxicity & Propaganda. We then follow the same methodology to obtain the CM and CMP image-generation prompt sets as above, only skipping conversion from direct to indirect prompts.
3.2 Models Evaluated
In this study, we benchmark four instruction-tuned LLMs of comparable sizes $\approx$ 8B parameters, with varying levels of multilingual capabilities33footnotemark: 3ChatGPT-4o-mini Hurst et al. (2024), Llama-3-8B-Instruct Dubey et al. (2024), Gemma-1.1-7b-it Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023). For the multimodal image generation task, we examine ChatGPT-4o-mini Hurst et al. (2024) and Gemini-2.5-Flash-Image Comanici et al. (2025).
3.3 Jailbreak Templates
In this study, we benchmark three existing jailbreak templates on the LLMs 3Opposite Mode (OM), AntiLM, AIM Shen et al. (2024b) across the English, CM and CMP input prompts. We extend the dual-persona concept of OM to simulate a resilience testing environment to create the Sandbox template. For the image generation task, we test with a Base template– instructing image generation without requesting clarifications on generation style. We also devise a new jailbreaking template– VisLM, which instructs the model to ‘forget’ its text generation capabilities and directly pass the text inputs to its image generator without any filtering. Both Base and VisLM templates are instructed to generate an image when the inputs are prefixed with ‘Input: ’.
4 Experimental Methodology
We describe the method of generating the code-mixed and phonetically perturbed prompts, followed by a brief overview of the evaluation metrics.
4.1 Generating Input Prompts
@@
We use a 3-step process to generate the English, CM and CMP prompt sets for both text and image generation tasks 3.
@@
    Questions $\rightarrow$ Hypothetical scenario Bhardwaj and Poria (2023): We convert input questions in the dataset to hypothetical scenario statements, obtaining the English set.
@@ 4.2 Evaluation Metrics
We evaluate the outputs of both our text and image generation tasks using the metrics described as follows. An input to a model is a four-tuple that generates a response– $R=\langle M,J,P,T\rangle$, where the model is $M$, jailbreak template $J$, the prompt (English/CM/CMP) is $P$, and temperature is $T\in\{0.2k\mid k=0,1,2,3,4,5\}$. For LLMs, we evaluate for all temperature values and report the average. We do not experiment with multiple temp. values on MLLMs due to feature unavailability and financial constraints.
Success & Relevance: We use GPT-4o-mini as an LLM-as-a-judge Zheng et al. (2023) to quantify the success and relevance of the generated responses. A binary function, $\mathbb{S}(R)$, returns ‘1’ if the attack is successful and ‘0’ otherwise. Similarly, a ternary function $\mathbb{R}(R)$ returns ‘1’ if the attack is relevant, ‘0’ if irrelevant and ‘-1’ for refusal. To evaluate the effectiveness of our automatic judge, we conduct a binary annotation experiment wherein 3 volunteers annotate 100 outputs generated by ChatGPT in the ‘None’ case on the English set across the 6 temperature values. We report the ICC Bartko (1966) between the human and GPT judge scores.
@@
4.3 Interpreting Phonetic Perturbations
To understand how phonetic perturbations successfully bypass the safety filters of models, we conduct an interpretability experiment on Llama-3-8B-Instruct. We describe the methodology as follows. (1) We select a small subset of the dataset, specifically with $ASR_{CM}\leq 0.33, AA$$ASR_{CMP}\geq 0.5$ while ensuring that $AARR_{CMP}\geq AARR_{CM}$.
@@ 4.3 Interpreting Phonetic Perturbations
(3) For prompts in all three formats– English, CM and CMP, we use LayerIntegratedGradients, Captum’s LLM variant for Integrated Gradients Sundararajan et al. (2017) to generate sequence attribution bar plots– token-wise attribution (importance) scores for the generation of a safe response from the model. In each plot, we discard the tokens with an attribution score $S\in[-0.20,0.20]$.

### v4

# Code-Mixed Phonetic Perturbations for Red-Teaming LLMs

@@ Abstract
Large language models (LLMs) continue to be demonstrably unsafe despite sophisticated safety alignment techniques and multilingual red-teaming. However, recent red-teaming work has focused on incremental gains in attack success over identifying underlying architectural vulnerabilities in models. In this work, we present CMP-RT, a novel red-teaming probe that combines code-mixing with phonetic perturbations (CMP), exposing a tokenizer-level safety vulnerability in transformers. Combining realistic elements from digital communication such as code-mixing and textese, CMP-RT preserves phonetics while perturbing safety-critical tokens, allowing harmful prompts to bypass alignment mechanisms while maintaining high prompt interpretability, exposing a gap between pre-training and safety alignment. Our results demonstrate robustness against standard defenses, attack scalability, and generalization of the vulnerability across modalities and to SOTA models like Gemini-3-Pro, establishing CMP-RT as a major threat model and highlighting tokenization as an under-examined vulnerability in current safety pipelines.
@@
Code-Mixed Phonetic Perturbations for Red-Teaming LLMs
@@ 1 Introduction
The wide-scale deployment of large language models (LLMs) across both general-purpose Hadi et al. (2023) and safety-critical Hua et al. (2024) tasks has led to increased scrutiny on their safety Salhab et al. (2024). Red teaming Sarkar (2025) uses prompting strategies Pang et al. (2025) to bypass safety filters of LLMs and elicit harmful or unethical responses Wei et al. (2023) that exposes model biases and vulnerabilities. Recent red-teaming work Hughes et al. (2024); Li et al. (2024a) has emphasized optimization-driven jailbreaking that maximize attack success rates using complex, often uninterpretable inputs. While effective, such approaches often provide limited insight into why safety mechanisms fail. In contrast, we study red-teaming as a diagnostic process, introducing a multilingual probe that isolates a concrete failure mode at the input level representation, enabling direct analysis of the underlying vulnerability.
A particularly challenging area of safety research is multilingual alignment Wang et al. (2024b). LLM pre-training data exposes models to various forms of digital communication phenomena such as code-switching Gardner-Chloros (2009)– mixing languages using their original scripts (common in verbal communication Li et al. (2024b), and code-mixing Thara and Poornachandran (2018)– mixing languages using a single, primary script (widely observed on online platforms and SMS-conversations Das and Gambäck (2013)). Moreover, English speakers in non-Anglophone societies often create new, possibly strange spellings for words based on their phonetic perceptions (‘design’ $\rightarrow$ ‘dezain’). This is often observed in textese Drouin (2011), a form of communication common in SMS and internet conversations Thakur (2021), manifesting as informal, correct-sounding misspellings (phonetic perturbations) in LLM pre-training data. However, alignment strategies typically utilize standardized multilingual Ropers et al. (2024), code-switched Yoo et al. (2024) or code-mixed Bohra et al. (2018) inputs that preserve canonical spellings and are easier to generate and control programmatically. This creates a representational gap between the informal perturbations seen during pre-training and the inputs used to train or evaluate safety mechanisms.
This work presents CMP-RT, a novel red-teaming probe combining code-mixing (CM) with phonetic perturbations (CMP) – injecting textese-style phonetically similar sounding spelling errors in sensitive words – exposing a major safety vulnerability at the tokenizer-level. Our study poses the following research questions.
@@
RQ1. Do the safety guardrails of SOTA LLMs generalise to textese-style code-mixed (CMP) inputs? While models undergo extensive multilingual alignment Dubey et al. (2024), real-world informal interactions [^1] https://www.bcg.com/publications/2024/consumers-know-more-about-ai-than-businesses-think necessitate evaluation of such cases. Our results indicate that CMP increases harmful outputs, revealing the lack of generalisability of existing safeguards.
RQ2. Does CMP-RT trigger relevant responses from the LLMs? CMP-RT intentionally misspells words while preserving the pronunciation, necessitating an input interpretability analysis. We show that LLMs generate harmful outputs despite demonstrating high input interpretability, revealing a wide gap between pre-training & safety.
@@
RQ3. What failure mechanism does CMP-RT exploit to bypass LLM safety filters? To identify the actual vulnerability exploited by CMP-RT, we utilize Integrated Gradients Sundararajan et al. (2017) to conduct an interpretability analysis on an open-source LLM, revealing that CMP-RT alters word tokenization in a way that suppresses attributions from safety-critical inputs tokens.
RQ4. How generalizable and scalable is the vulnerability exposed by CMP-RT? We investigate the effectiveness of CMP-RT on Multimodal LLMs (MLLMs) and the scalability of the approach through supervised finetuning, demonstrating that our approach is both highly generalizable to MLLMs as well as scalable using high-quality, hand-crafted data as seed.
@@ 2 Related Work
Red-Teaming: Red teaming Ganguli et al. (2022) focuses on evaluating LLMs for safety and vulnerability concerns Bhardwaj and Poria (2023). Jailbreaking is one such method that involves bypassing the safety training of LLMs to elicit harmful or unethical outputs. While white-box jailbreak techniques require access to model weights for attack optimization Wang et al. (2024a), black-box methods Mehrotra et al. (2024) rely on prompting techniques to probe models and hence are not restricted to open-source models. Recent work, in addition to text Chen et al. (2025); Liu et al. (2023), has extended evaluations to multiple modalities Liu et al. (2024); Song et al. (2025) and languages Ropers et al. (2024).
Code-Mixing: Code-mixing (CM) – a special form of multilingualism that combines multiple languages using a primary script – has helped increase the performance Shankar et al. (2024) and capabilities Zhang et al. (2024) of LLMs in multilingual settings. Prior multilingual safety work involves comprehensive evaluations in multiple languages Shen et al. (2024a), alignment Song et al. (2024) strategies, and even red-teaming model in code-switched Yoo et al. (2024) settings. However, code-mixing demonstrates a special case where only one of the languages is in it’s original script, closely resembling digital communication Thara and Poornachandran (2018). Moreover, in such informal settings, users (specially from non-Anglophone societies) often depart from canonical spellings through pronunciation-preserving misspellings, known as textese Drouin (2011); Thakur (2021). Despite their prevalence, these textese-style variations are rarely considered in multilingual safety evaluations, mandating the evaluation of models under such realistic input conditions.
In this study, we study code-mixing with phonetic perturbations (altering word spellings while preserving pronunciation and semantic meaning) as a red-teaming probe to uncover a major architectural safety vulnerability in LLMs. Our jailbreak strategy successfully jailbreaks SOTA models like Llama 3, ChatGPT 4o-mini and Gemini 3 Pro for both text and image generation tasks.
3 Experimental Setup
3.1 Code-Mixing with Phonetic Perturbations
@@
We use a 3-step process to generate the English, CM and CMP prompt sets for both text and image generation tasks 2. We use the Hindi language as our medium of code-mixing.
@@
    Questions $\rightarrow$ Hypothetical scenario Bhardwaj and Poria (2023): We convert the default inputs in the dataset (Default-set) to hypothetical scenarios, obtaining the English set.
@@
3.2 Datasets Benchmarked
Prompts for text generation: We utilize three benchmark datasets 2HarmfulQA Bhardwaj and Poria (2023), NicheHazardQA Hazra et al. (2024) and TechHazardQA Banerjee et al. (2024) – studying model vulnerabilities, refusal training and compliance with harmful queries. For a comprehensive evaluation of our red-teaming strategy, we sample 20 prompts from each category in each dataset, yielding a total of 460 prompts across 23 categories. All prompts are originally in the English language. Thus, using the sampled datasets, we manually generate the CM and CMP prompt sets, described later.
Prompts for image generation: Using a set of 10 handwritten samples, we prompt GPT-4o to automatically generate the image generation Default-sets of 20 red-teaming prompts each to test the model’s resilience against various categories of harm – Religious Hate, Casteist Hate, Gore, Self-Harm and Social Media Toxicity & Propaganda. We then follow the same methodology to obtain the CM and CMP image-generation prompt sets as above, only skipping conversion from direct to indirect prompts.
3.3 Models Evaluated
Text generation models: We benchmark four instruction-tuned LLMs of comparable sizes $\approx$ 8B parameters, with varying levels of multilingual capabilities 2ChatGPT-4o-mini Hurst et al. (2024), Llama-3-8B-Instruct Dubey et al. (2024), Gemma-1.1-7b-it Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023).
Image generation models: We benchmark ChatGPT-4o-mini Hurst et al. (2024), Gemini-2.5-Flash-Image and Nano Banana Pro based on Gemini-3-Pro Comanici et al. (2025).
3.4 Jailbreak Templates
Templates for text generation: We benchmark using four jailbreak templates, three existing 2Opposite Mode (OM), AntiLM, AIM Shen et al. (2024b) across the English, CM and CMP input prompts. We extend the dual-persona concept of OM to simulate a resilience testing environment to create the fourth– Sandbox template.
Templates for image generation: For the image generation task, we test with a Base template– instructing image generation without requesting clarifications on generation style. We also devise a new jailbreaking template– VisLM, which instructs the model to ‘forget’ its text generation capabilities and directly pass the text inputs to its image generator without any filtering. Both Base and VisLM templates are instructed to generate an image when the inputs are prefixed with ‘Input: ’.
@@ 3.5 Evaluation Metrics
We evaluate the outputs of both our text and image generation tasks using the metrics described as follows. An input to a model is a four-tuple that generates a response– $R=\langle M,J,P,T\rangle$, where the model is $M$, jailbreak template $J$, the prompt (English/CM/CMP) is $P$, and temperature is $T\in\{0.2k\mid k=0,1,2,3,4,5\}$. For LLMs, we evaluate for all temperature values and report the average. We do not experiment with multiple temperature values on MLLMs due to feature unavailability and financial constraints.
Success & Relevance: We use GPT-4o-mini as an LLM-as-a-judge Zheng et al. (2023) to quantify the success and relevance of the generated responses. A binary function, $\mathbb{S}(R)$, returns ‘1’ if the attack is successful and ‘0’ otherwise. Similarly, a ternary function $\mathbb{R}(R)$ returns ‘1’ if the response is relevant, ‘0’ if irrelevant and ‘-1’ for refusal.
LLM-Judge Validation: We conduct a binary annotation experiment wherein 3 volunteers annotate the outputs generated by ChatGPT on the English set and Gemma on the CMP set for 100 prompts across the 6 temperature values, both for the ‘None’ case. We report the ICC Bartko (1966) between the human and GPT judge scores.
@@
3.6 Experiments
3.6.1 Red-Teaming LLMs with CMP-RT
First, we evaluate the effectiveness of CMP-RT against the standard English and CM attacks across our model-set, jailbreak templates, and temperature settings. For each configuration, we measure AASR– capturing the ability of the input to elicit harmful outputs, and AARR– quantifying whether the generated responses remain aligned with the original prompt intent.
3.6.2 Robustness against Adversarial Defenses
To verify the non-triviality of CMP-RT and establish the severity of the threat model, we test our attack against the following defenses.
OpenAI Moderation API: We compare the average flag-rates of the CMP and English sets.
Perplexity-based Filtering Alon and Kamfonas (2023): We finetune GPT-4o-mini on our CMP set to automatically convert 460 general purpose prompts sourced from Databricks Dolly 15k  Conover et al. (2023) (Safe-Default) into Safe- English, CMP sets. We compare the average GPT-2 perplexities for each safe-harmful pair.
3.6.3 Benchmarking CMP-RT
We compare our attack strategy against CSRT Yoo et al. (2024), a multilingual red-teaming strategy which utilizes code-switching in multiple languages to jailbreak LLMs. We convert our Default and the English sets into CSRT-style inputs 2– Default-CSRT (directly corresponds to the original CSRT attack) and Hypothetical-CSRT, evaluating them against the CMP set across the same set of models and temperature settings using the ‘None’ template and reporting the AASR and AARR.
3.6.4 Interpreting Phonetic Perturbations
Next, to reveal the failure mechanism exploited by CMP-RT, we conduct an interpretability experiment on Llama-3-8B-Instruct. We describe the methodology as follows.
(1) We select a small subset of the dataset, specifically with $ASR_{CM}\leq 0.33, AA$$ASR_{CMP}\geq 0.5$ while ensuring that $AARR_{CMP}\geq AARR_{CM}$.
@@ 3.6.4 Interpreting Phonetic Perturbations
(3) For prompts in all three formats– English, CM and CMP, we use LayerIntegratedGradients from Captum Sundararajan et al. (2017) to generate sequence attribution bar plots– token-wise attribution (importance) scores for the generation of a safe response from the model. In each plot, we discard the tokens with an attribution score $S\in[-0.20,0.20]$.
@@
3.6.5 Multimodal and Automated CMP-RT
Next, we examine the generalization of CMP-RT to the image modality. We evaluate our image-generation English, CM and CMP sets, reporting AASR and AARR for the generated images across the configurations previously reported.
Lastly, we fine-tune three variants of Llama-3-8B-Instruct to automate the successive stages of the CMP generation pipeline: Default prompt $\rightarrow$ (hypothetical) English prompt $\rightarrow$ CM prompt $\rightarrow$ CMP prompt. Next, we automatically convert 521 prompts sourced from AdvBench Zou et al. (2023) into a test CMP set using the pipeline and report AASR and AARR evaluated on Llama-3-8B-Instruct at a fixed temperature value of 0.5.
Thus, for the text-generation experiment, each LLM is evaluated across 5 jailbreak templates, 3 prompt sets consisting of 460 prompts each, and 6 temperature values, resulting in a total of 41,400 responses. For the image generation experiment, each model is evaluated across 2 jailbreak templates and 3 prompt sets, each consisting of 110 prompts, generating a total of 660 responses.